### v2

@@ Abstract
To predict high-resolution (HR) omnidirectional depth map, existing methods typically leverage HR omnidirectional image (ODI) as the input via fully-supervised learning. However, in practice, taking HR ODI as input is undesired due to resource-constrained devices. In addition, depth maps are often with lower resolution than color images. Therefore, in this paper, we explore for the first time to estimate the HR omnidirectional depth directly from a low-resolution (LR) ODI, when no HR depth GT map is available. Our key idea is to transfer the scene structural knowledge from the HR image modality and the corresponding LR depth maps to achieve the goal of HR depth estimation without any extra inference cost. Specifically, we introduce ODI super-resolution (SR) as an auxiliary task and train both tasks collaboratively in a weakly supervised manner to boost the performance of HR depth estimation. The ODI SR task extracts the scene structural knowledge via uncertainty estimation. Buttressed by this, a scene structural knowledge transfer (SSKT) module is proposed with two key components. First, we employ a cylindrical implicit interpolation function (CIIF) to learn cylindrical neural interpolation weights for feature up-sampling and share the parameters of CIIFs between the two tasks. Then, we propose a feature distillation (FD) loss that provides extra structural regularization to help the HR depth estimation task learn more scene structural knowledge. Extensive experiments demonstrate that our weakly-supervised method outperforms baseline methods, and even achieves comparable performance with the fully-supervised methods.
@@
Our weakly-supervised can adapt to various backbones, input sizes, and datasets. Our method can also easily be applied to higher resolution, such as $1024\times 2048$. We contribute the effectiveness to the proposed SSKT module, which successfully transfers essential scene structural knowledge from the ODI SR task to HR depth estimation task. Importantly, we prove that the way of transferring the scene structural knowledge from the HR image modality and corresponding LR depth map is feasible to achieve the goal of HR depth estimation without any extra inference cost.
@@ 1 Introduction
Monocular depth estimation is a fundamental task of 3D scene understanding that finds its applications in various fields, such as autonomous driving. The advancements in deep neural networks (DNNs) [1] and the accessibility of large-scale annotated datasets have resulted in notable improvements in monocular depth estimation. Nevertheless, most of the existing approaches are developed for perspective images, which suffer from a limited field-of-view (FoV) [2, 3, 4, 5, 6, 7, 8].
@@
Recently, $360^{\circ}$ cameras [^1] Also called omnidirectional cameras. have received considerable attention because of their ability to capture the environment with a wide FoV of $180^{\circ}\times 360^{\circ}$ in a single shot. Omnidirectional images (ODIs) [^2] ODIs usually indicate the equirectangular projection (ERP) type images. present the scene entirely with rich scene contexts. This distinct advantage has inspired active research for monocular depth estimation [9, 10, 11, 12]. However, the angular resolution of ODIs is relatively lower than that of perspective images when ODIs and perspective images are captured by different sensors with the identical size. The low angular resolution leads to degraded structural details such as edges, rendering it difficult for omnidirectional depth estimation.
@@
To reconstruct the structural details, existing methods [12, 10, 13, 11, 14, 15, 16, 17, 18] leverage high-resolution (HR) ODIs as inputs with the supervision of HR depth ground truth (GT) maps. For instance, using the HR ODIs and paired HR depth GT maps, OmniDepth [14] employs the encoder-decoder architecture to predict HR depth maps, as shown in Fig. 1(a). However, in practice, taking HR ODI as input is undesired by resource-constrained devices. Take virtual reality (VR) as an example, the headsets often have limited bandwidth for transmission and limited computational resources [19, 20]. Processing HR ODIs in the headsets might cause high latency due to limited bandwidth [19]. In addition, obtaining HR depth maps is arduous and expensive [21]. This is because depth sensors, such as LiDARs, often capture sparse depths with low resolution [22]. Also, increasing the lens of LiDARs is far more expensive than obtaining HR ODI [23].
@@
Motivation. In this paper, we explore for the first time to estimate the HR depth map directly from an LR ODI, when no HR depth GT map is available (See Fig 1(b)). Without the supervision of HR depth GT maps, the fine scene structural details can scarcely be recovered in the HR depth predictions. Therefore, to explore valuable supplementary supervision that can facilitate HR depth estimation, our key idea is to learn and transfer scene structural knowledge from the HR image modality and the corresponding LR depth maps to achieve the goal of HR depth estimation without causing extra inference cost. This is inspired by the fact that 1) there exist publicly available large-scale LR-HR paired ODIdataset and thus achieve the ODI super-resolution (SR) task [24, 25]; 2) an ODI and its corresponding depth map are photometric and geometric representations of the same scene, respectively [21]. This implies that both representations share the common scene structural knowledge, such as edges [26].
@@
Firstly, considering the unique properties of the equirectangular projection (ERP) type of ODIs, we design a novel cylindrical implicit interpolation function (CIIF) for feature up-sampling by learning the neural interpolation weights based on cylindrical coordinates. This way, our CIIF alleviates the discontinuity and stretched distortion of ODIs. Also, as the learned interpolation weights can capture the correlations among neighboring pixels, they can be used to represent the scene structural knowledge. Therefore, by sharing the parameters of CIIFs between the ODI SR and HR depth estimation tasks, structural knowledge can be transferred between the two tasks effectively. Secondly, as the features of the ODI SR task also contain crucial scene structural knowledge, we propose a feature distillation (FD) loss to further boost the HR depth estimation performance.
In total, our weakly-supervised method is efficient, as the ODI SR network and SSKT module can be removed freely and only LR ODI is needed during inference, introducing no extra inference cost. Moreover, our method is flexible and can be easily adapted as a plug-and-play approach by modifying the backbones for the HR depth estimation network and varying the up-sampling factors (See Tab. 1). Furthermore, we demonstrate the effectiveness of our method across different datasets, [29, 30, 14], with various up-sampling factors and backbones. The experimental results show that our method not only outperforms the baseline methods, but also achieves comparable performance with the fully-supervised methods. Our main contributions can be summarized into four-fold:
  • We propose the first weakly-supervised method that predicts the HR omnidirectional depth map from an LR ODI, when no HR depth GT map is available.
@@
  • We propose an SSKT module with two components. The first is the CIIF as it is specific for ODIs to learn neural interpolation weights, whose parameters are shared between the two tasks. The second is the FD loss to provide extra structural regularization.
  • Experimental results demonstrate that our weakly supervised method is efficient, flexible, and effective, which can achieve comparable results with the fully supervised methods.
@@ 2 Related Work
Omnidirectional Monocular Depth Estimation. Some attempts have been made to tackle the spherical distortion in ODIs by modifying the characteristics of convolutional filters, including row-wise rectangular filter [14], deforming the sampling grids [16], and dilated convolutions for larger receptive fields [17]. Another direction is to combine different projection types, e.g., ERP, cube map [12, 10], and tangent projection [13, 31]. BiFuse [12] proposes a bi-projection fusion model for the ERP images and cube maps to share information between each other, while UniFuse [10] demonstrates that the unidirectional fusion from the cube map to ERP is more effective. Some methods [15, 18] focus on parallel computation in omnidirectional monocular depth estimation. The typical pipeline is first projecting an ERP image into multiple low-resolution perspective patches, then predicting depth maps in parallel, and finally re-projecting the perspective depth predictions back to the ERP plane. Recently, Li et al. [32] propose to extract features on the unit sphere to relieve the distortion effect. However, all the aforementioned methods require the HR depth maps for supervision.
@@
Omnidirectional image super-resolution. Early ODI SR methods [33, 34, 35] assemble multiple LR ODIs to reconstruct an HR one. However, they rely heavily on the precision of calibration and the number of inputs. Cagri et al. [36] propose the first DNN-based framework via adversarial learning. LAU-Net [24] addresses the non-uniformly distributed pixel density of the ERP type ODI via adaptively up-sampling different latitude bands. SphereSR [25] proposes a continuous spherical representation to model the spherical data as an icosahedron and reconstruct ODIs with multiple projection types. Recently, OSRT [37] and OPDN [38] employ distortion-aware transformers to alleviate the effect of distortion in ERP type ODI.
@@
Modeling Uncertainty in image super-resolution. In the image SR task, high-frequency regions with rich texture and edges carry more visual information than the smooth regions [28]. To better capture the high-frequency details and emphasize the corresponding pixels, some uncertainty-driven image SR methods have been proposed [39, 28]. For instance, GRAM [39] extracts the uncertainty map via the Gaussian assumption and takes the uncertainty map as the attenuation mask to achieve better performance. Differently, Ning et al. [28] estimate the uncertainty with Jeffrey’s prior [40] and proposes an uncertainty-driven loss to provide the attention for the pixels with high uncertainty. Our method introduces the uncertainty estimation to extract scene structural knowledge, which is transferred to depth estimation task.
@@
Implicit Neural Representation. It uses few parameters with the help of multi-layer perceptrons (MLPs) [41, 42, 43]. For the image SR task, Chen et al. [44] propose an implicit function, LIIF, to represent images continuously and achieves the up-sampling process with arbitrary factor. Furthermore, Tang et al. [45] replace the non-parametric interpolation weights in LIIF with the ones learned from the input image and depth map to achieve the guided depth SR task. However, the two mentioned implicit functions are based on the Cartesian coordinate system, which is weak in tackling the discontinuity and distortion problems in the ERP type ODIs. In contrast, our proposed cylindrical implicit interpolation function (CIIF) is constructed based on cylindrical coordinates, which is effective to alleviate the mentioned problems in ERP type ODIs.
@@
Multi-task learning. It is mainly used for correlated tasks [46], e.g., semantic segmentation and depth estimation [47]. Liu et al. [48] leverage 3D features to enhance 2D features extracted from the images. Wang et al. [49] treat image SR as an auxiliary task for semantic segmentation with fully supervision. Sun et al. [21] leverage the monocular depth estimation task to facilitate the depth SR task, reconstructing finer the structural details. In addition, Yang et al. [50] proposes a cross-domain attention conciliation module to enhance the depth feature when encoding the image and depth features. Different from these methods, based on the geometric similarity, we aim to transfer structural knowledge from the fully-supervised ODI SR task to the weakly-supervised HR depth estimation task.
@@ 3 Method
Overview. The goal is to predict an HR omnidirectional depth map $D_{\mathrm{HR}}$ from an LR ODI $I_{\mathrm{LR}}$ as input, using only an LR depth map $D^{\mathrm{GT}}_{\mathrm{LR}}$ for supervision. As shown in Fig. 2, our framework consists of an ODI SR task, a scene structural knowledge transfer (SSKT) module, and an HR depth estimation task. The ODI SR is employed as an auxiliary task and trained collaboratively with the HR depth estimation task. Specifically, we introduce uncertainty estimation to the ODI SR task for scene structural knowledge extraction. To transfer the extracted knowledge, the SSKT module is proposed with two components. The first is to share the parameters of the proposed cylindrical implicit interpolation function (CIIF) between the two tasks. The second is the feature distillation (FD) loss that provides extra structural regularization. We now describe the components in detail.
@@ 3.1 ODI SR
The omnidirectional images contain not only the complete structural details of the surroundings, but also rich textural details in the scenes. However, the enriched textural details can interfere with the smoothness of predicted depth maps [11, 10, 27]. Therefore, directly transferring knowledge from the ODI SR task degrades the depth estimation performance (See Tab. 3). Inspired by the recent progress of uncertainty-driven image super-resolution [28], which demonstrates that regions with high uncertainty are often matched with high-frequency regions, we introduce the uncertainty estimation to the ODI SR task. By doing so, the regions that carry more structural knowledge can be emphasized, which is beneficial to reconstruct more structural details in the HR depth estimation task. Specifically, as shown in Fig. 2, we utilize the feature extractor to process the LR ODI $I_{\mathrm{LR}}$, and utilize two groups of convolutional layers with ELU activation functions [51] to output LR features for the RGB part $F_{\mathrm{LR}}^{\mathrm{RGB}}$ and uncertainty part $F_{\mathrm{LR}}^{\mathrm{U}}$, respectively. Then, $F_{\mathrm{LR}}^{\mathrm{RGB}}$ and $F_{\mathrm{LR}}^{\mathrm{U}}$ are up-sampled by CIIFs with parameters $g_{\rho}$ and $g_{\theta}$ to generate HR features $F_{\mathrm{HR}}^{\mathrm{RGB}}$ and $F_{\mathrm{HR}}^{\mathrm{U}}$, respectively. We then describe the components of CIIF that achieve the feature up-sampling process.
CIIF. As depicted in Fig. 3(a), the image interpolation problems can be defined as follows:
@@
where $x_{q}$ is the coordinate of query pixel $q$ in the HR image domain, $x_{i}$ is the coordinate of corner pixel $i$ in the LR image domain; $N_{q}$ represents the neighboring pixels of $q$ in the LR image domain and $w_{q,i}$ is the interpolation weight between pixel $q$ and $i$. In LIIF [44], as shown in Fig. 3(a), given the latent code $z_{i}$ encoded from $I_{\mathrm{LR}}$ and the coordinate relation between $q$ and $i$, the RGB value of $q$ in the HR image domain is predicted by the MLP $f_{\theta}$ with parameters $\theta$ as follows:
@@
Furthermore, [45] replaces the scale factor $S_{i}$ in Eq. (3) with learnable interpolation weights to obtain smoother results. However, for ODIs with ERP type, which are unfold from the sphere in a cylindrical manner, there are two specific challenges: 1) The discontinuity between the left and right sides; 2) The stretched distortion toward the poles (the top and bottom of ODIs). Previous implicit interpolation methods, e.g[44, 45], are based on the Cartesian coordinate, which are less effective in representing spatial correlation of pixels in ODIs [25]. Therefore, we propose to reframe ODIs with the cylindrical representation, as shown in Fig. 3(b). The cylindrical angle coordinate is formulated as $\alpha\in(-\arctan(\frac{\pi h}{w}),\ \arctan(\frac{\pi h}{w}))$ (Latitude angle), $\beta\in(-\pi,\ \pi)$ (Longitude angle). For the calculation of the cylindrical representation, we recommend readers to refer to the first section of the supplemental material. Based on the cylindrical angle coordinate, we learn the cylindrical implicit interpolation weights (See Fig. 3(b)):
@@
where $N$ is the number of pixels, and $I^{\mathrm{GT}}_{\mathrm{HR}}$ is HR GT image.

### v3

@@ Abstract
To predict high-resolution (HR) omnidirectional depth maps, existing methods typically leverage HR omnidirectional image (ODI) as the input via fully-supervised learning. However, in practice, taking HR ODI as input is undesired due to resource-constrained devices. In addition, depth maps are often with lower resolution than color images. Therefore, in this paper, we explore for the first time to estimate the HR omnidirectional depth directly from a low-resolution (LR) ODI, when no HR depth GT map is available. Our key idea is to transfer the scene structural knowledge from the HR image modality and the corresponding LR depth maps to achieve the goal of HR depth estimation without any extra inference cost. Specifically, we introduce ODI super-resolution (SR) as an auxiliary task and train both tasks collaboratively in a weakly supervised manner to boost the performance of HR depth estimation. The ODI SR task extracts the scene structural knowledge via uncertainty estimation. Buttressed by this, a scene structural knowledge transfer (SSKT) module is proposed with two key components. First, we employ a cylindrical implicit interpolation function (CIIF) to learn cylindrical neural interpolation weights for feature up-sampling and share the parameters of CIIFs between the two tasks. Then, we propose a feature distillation (FD) loss that provides extra structural regularization to help the HR depth estimation task learn more scene structural knowledge. Extensive experiments demonstrate that our weakly-supervised method outperforms baseline methods, and even achieves comparable performance with the fully-supervised methods.
Abstract
We provide more details of the proposed method in the supplementary material. Specifically, in Sec. .1 we explain explicitly the proposed cylindrical implicit interpolation function (CIIF) and its sharing strategy. Sec. .2 shows more details in the feature distillation (FD) loss. Then, Sec. .3 shows the more experimental results, including datasets and metrics, visual comparisons, ablation studies, and reports on inference time.
@@
Our weakly-supervised can adapt to various backbones, input sizes, and datasets. Our method can also easily be applied to higher resolution, such as $1024\times 2048$. We contribute the effectiveness to the proposed SSKT module, which successfully transfers essential scene structural knowledge from the ODI SR task to the HR depth estimation task. Importantly, we prove that the way of transferring the scene structural knowledge from the HR image modality and corresponding LR depth map is feasible to achieve the goal of HR depth estimation without any extra inference cost.
@@ 1 Introduction
Monocular depth estimation is a fundamental task of 3D scene understanding that finds its applications in various fields, such as autonomous driving. The advancements in deep neural networks (DNNs) [1] and the accessibility of large-scale annotated datasets have resulted in notable improvements in monocular depth estimation. Nevertheless, most of the existing approaches are developed for perspective images, which suffer from a limited field-of-view (FoV) [2, 3].
@@
Recently, $360^{\circ}$ cameras [^1] Also called omnidirectional cameras. have received considerable attention because of their ability to capture the environment with a wide FoV of $180^{\circ}\times 360^{\circ}$ in a single shot. Omnidirectional images (ODIs) [^2] ODIs usually indicate the equirectangular projection (ERP) type images. present the scene entirely with rich scene contexts. This distinct advantage has inspired active research for monocular depth estimation [4, 5, 6, 7]. However, the angular resolution of ODIs is relatively lower than that of perspective images when ODIs and perspective images are captured by different sensors with identical sizes. The low angular resolution leads to degraded structural details such as edges, rendering it difficult for omnidirectional depth estimation.
@@
To reconstruct the structural details, existing methods [7, 5, 8, 6, 9, 10, 11, 12, 13] leverage high-resolution (HR) ODIs as inputs with the supervision of HR depth ground truth (GT) maps. For instance, using the HR ODIs and paired HR depth GT maps, OmniDepth [9] employs the encoder-decoder architecture to predict HR depth maps, as shown in Fig. 1(a). However, in practice, taking HR ODI as input is undesired by resource-constrained devices. Take virtual reality (VR) as an example, the headsets often have limited bandwidth for transmission and limited computational resources [14]. Processing HR ODIs in the headsets might cause high latency due to limited bandwidth [14]. In addition, obtaining HR depth maps is arduous and expensive [15]. This is because depth sensors, such as LiDARs, often capture sparse depths with low resolution [16]. Also, increasing the lens of LiDARs is far more expensive than obtaining HR ODI [17].
@@
Motivation. In this paper, we explore for the first time to estimate the HR depth map directly from an LR ODI, when no HR depth GT map is available (See Fig 1(b)). Without the supervision of HR depth GT maps, the fine scene structural details can scarcely be recovered in the HR depth predictions. Therefore, to explore valuable supplementary supervision that can facilitate HR depth estimation, our key idea is to learn and transfer scene structural knowledge from the HR image modality and the corresponding LR depth maps to achieve the goal of HR depth estimation without causing extra inference cost. This is inspired by the fact that 1) there exists publicly available large-scale LR-HR paired image datasets and thus achieve the super-resolution (SR) task [18, 19, 20, 21, 22]; 2) an ODI and its corresponding depth map are photometric and geometric representations of the same scene, respectively [15]. This implies that both representations share the common scene structural knowledge, such as edges.
@@
Firstly, considering the unique properties of the equirectangular projection (ERP) type of ODIs, we design a novel cylindrical implicit interpolation function (CIIF) for feature up-sampling by learning the neural interpolation weights based on cylindrical coordinates. This way, our CIIF alleviates the discontinuity and stretched distortion of ODIs. Also, as the learned interpolation weights can capture the correlations among neighboring pixels, they can be used to represent the scene structural knowledge. Therefore, by sharing the parameters of CIIFs between the ODI SR and HR depth estimation tasks, structural knowledge can be transferred between the two tasks effectively. Secondly, as the features of the ODI SR task also contain crucial scene structural knowledge, we propose a feature distillation (FD) loss to refine depth estimation.
In total, our weakly-supervised method is efficient, as the ODI SR network and SSKT module can be removed freely and only LR ODI is needed during inference, introducing no extra inference cost. Moreover, our method is flexible and can be easily adapted as a plug-and-play approach by modifying the backbones for the HR depth estimation network and varying the up-sampling factors (See Tab. 1). Furthermore, we demonstrate the effectiveness of our method across different datasets [25, 26, 9], with various up-sampling factors and backbones. The experimental results show that our method not only outperforms the baseline methods, but also achieves comparable performance with the fully-supervised methods. Our main contributions can be summarized into four-fold:
  • We propose the first weakly-supervised method that predicts the HR omnidirectional depth map from an LR ODI when no HR depth GT map is available.
@@
  • We propose an SSKT module with two components. The first is the CIIF as it is specific for ODIs to learn neural interpolation weights. The second is the FD loss to provide extra structural regularization.
  • Experimental results demonstrate that our proposed method is efficient, flexible, and effective, which achieves comparable results with the fully supervised methods.
@@ 2 Related Work
Omnidirectional Monocular Depth Estimation. Some attempts have been made to tackle the spherical distortion in ODIs by modifying the characteristics of convolutional filters, including row-wise rectangular filter [9], deforming the sampling grids [11], and dilated convolutions for larger receptive fields [12]. Another direction is to combine different projection types, e.g., ERP, cube map [7, 5], and tangent projection [8, 27]. BiFuse [7] proposes a bi-projection fusion model for the ERP images and cube maps to share information with each other, while UniFuse [5] demonstrates that the unidirectional fusion from the cube map to ERP is more effective. Some methods [10, 13] focus on parallel computation in omnidirectional monocular depth estimation. The typical pipeline is first projecting an ERP image into multiple low-resolution perspective patches, then predicting depth maps in parallel, and finally re-projecting the perspective depth predictions back to the ERP plane. Recently, Li et al. [28] propose to extract features on the unit sphere to relieve the distortion effect. However, all the aforementioned methods require HR depth maps for supervision.
@@
Omnidirectional image super-resolution. Cagri et al. [29] propose the first DNN-based framework via adversarial learning. LAU-Net [18] addresses the non-uniformly distributed pixel density of the ERP type ODI via adaptively up-sampling different latitude bands. SphereSR [19] proposes a continuous spherical representation to model the sphere as an icosahedron and reconstruct ODIs with multiple projection types.
@@
Modeling Uncertainty in image super-resolution. In the image SR task, high-frequency regions with rich texture and edges carry more visual information than the smooth regions [24]. To better capture the high-frequency details and emphasize the corresponding pixels, some uncertainty-driven image SR methods have been proposed [30, 24]. For instance, GRAM [30] extracts the uncertainty map via the Gaussian assumption and takes the uncertainty map as the attenuation mask to achieve better performance. Differently, Ning et al. [24] estimate the uncertainty with Jeffrey’s prior and proposes an uncertainty-driven loss to provide attention for the pixels with high uncertainty. Our method introduces the uncertainty estimation to extract scene structural knowledge, which is transferred to the depth estimation task.
@@
Implicit Neural Representation. For the image SR task, Chen et al. [31] propose an implicit function, LIIF, to represent images continuously and achieves the up-sampling process with arbitrary factors. Furthermore, Tang et al. [32] replace the non-parametric interpolation weights in LIIF with the ones learned from the input image and depth map to achieve the guided depth SR task. However, the mentioned two implicit functions [31, 32] are based on the Cartesian coordinate system, which is weak in tackling the discontinuity and distortion problems in the ERP type ODIs. In contrast, our proposed CIIF is constructed based on cylindrical coordinates, which is effective in alleviating the mentioned problems in ERP type ODIs, as evaluated in Tab. 4.
@@
Multi-task learning. It is mainly used for correlated tasks [33] or domains [34, 35, 36]. Wang et al. [37] treat image SR as an auxiliary task for semantic segmentation with full supervision. Sun et al. [15] leverage the monocular depth estimation task to facilitate the depth SR task, reconstructing finer details.
@@ 3 Method
Overview. The goal is to predict an HR omnidirectional depth map $D_{\mathrm{HR}}$ from an LR ODI $I_{\mathrm{LR}}$ as input, using only an LR depth map $D^{\mathrm{GT}}_{\mathrm{LR}}$ for supervision. As shown in Fig. 2, our framework consists of an ODI SR task, a scene structural knowledge transfer (SSKT) module, and an HR depth estimation task. The ODI SR is employed as an auxiliary task and trained collaboratively with the HR depth estimation task. Specifically, we introduce uncertainty estimation to the ODI SR task for scene structural knowledge extraction. To transfer the extracted knowledge, the SSKT module is proposed with two components. The first is to share the parameters of the proposed cylindrical implicit interpolation function (CIIF) between the two tasks. The second is the feature distillation (FD) loss that provides extra structural regularization. The overall training process can be seen in Algorithm. 1. We now describe the components in detail.
@@ 3.1 ODI SR
The omnidirectional images contain not only the complete structural details of the surroundings but also rich textural details in the scenes. However, the enriched textural details can interfere with the smoothness of predicted depth maps [6, 5, 23]. Therefore, directly transferring knowledge from the ODI SR task degrades the depth estimation performance (See Tab. 3). Inspired by the recent progress of uncertainty-driven image super-resolution [24], which demonstrates that regions with high uncertainty are often matched with high-frequency regions, we introduce the uncertainty estimation to the ODI SR task. By doing so, the regions that carry more structural knowledge can be emphasized, which is beneficial to reconstruct more structural details in the HR depth estimation task. Specifically, as shown in Fig. 2, we utilize the feature extractor to process the LR ODI $I_{\mathrm{LR}}$ and utilize two groups of convolutional layers with ELU activation functions [38] to output LR features for the RGB part $F_{\mathrm{LR}}^{\mathrm{RGB}}$ and uncertainty part $F_{\mathrm{LR}}^{\mathrm{U}}$, respectively. Then, $F_{\mathrm{LR}}^{\mathrm{RGB}}$ is up-sampled by CIIF with parameters $g_{\rho}$ to generate HR features $F_{\mathrm{HR}}^{\mathrm{RGB}}$. Similarly, $F_{\mathrm{LR}}^{\mathrm{U}}$ is up-sampled by CIIF with parameters $g_{\theta}$ to generate HR features $F_{\mathrm{HR}}^{\mathrm{U}}$. We then describe the components of CIIF that achieve the feature up-sampling process.
CIIF. Generally, the image interpolation problems can be defined as follows:
@@
where $x_{q}$ is the coordinate of query pixel $q$ in the HR image domain, $x_{i}$ is the coordinate of corner pixel $i$ in the LR image domain; $N_{q}$ represents the neighboring pixels of $q$ in the LR image domain and $w_{q,i}$ is the interpolation weight between pixel $q$ and $i$. We take LIIF [31] as an example to illustrate the interpolation process, depicted in Fig. 3(a). In LIIF, given the latent code $z_{i}$ encoded from $I_{\mathrm{LR}}$ and the coordinate relation between $q$ and $i$, the RGB value of $q$ in the HR image domain is predicted by the MLP $f_{\theta}$ with parameters $\theta$ as follows:
@@
Furthermore, [32] replaces the scale factor $S_{i}$ in Eq. (3) with learnable interpolation weights to obtain smoother results. However, for ODIs with ERP type, which unfolds from the sphere in a cylindrical manner, there are two specific challenges: 1) The discontinuity between the left and right sides; 2) The stretched distortion toward the poles (the top and bottom of ODIs). Previous implicit interpolation methods, e.g[31, 32], are based on the Cartesian coordinate, which is less effective in representing spatial correlation of pixels in ODIs [19]. Therefore, we propose to reframe ODIs with the cylindrical representation, as shown in Fig. 3(b). The cylindrical angle coordinate is formulated as $\alpha\in(-\arctan(\frac{\pi h}{w}),\ \arctan(\frac{\pi h}{w}))$ (Latitude angle), $\beta\in(-\pi,\ \pi)$ (Longitude angle). For the calculation of the cylindrical representation, we recommend readers to refer to the first section of the supplemental material. Based on the cylindrical angle coordinate, we learn the cylindrical implicit interpolation weights (See Fig. 3(b)):
@@
Algorithm 1 The overall training process of our framework
1:  Training data: $I_{\mathrm{LR}},I_{\mathrm{HR}}^{\mathrm{GT}},D_{\mathrm{LR}}^{\mathrm{GT}}$
2:  for for i=1, $i\leq 30$ do
3:     ODI SR task:
4:     LR features $F_{\mathrm{LR}}^{\mathrm{RGB}}$ and $F_{\mathrm{LR}}^{\mathrm{U}}$ from $I_{\mathrm{LR}}$
5:     Up-sample $F_{\mathrm{LR}}^{\mathrm{RGB}}$ and $F_{\mathrm{LR}}^{\mathrm{U}}$ with CIIFs of parameters $g_{\rho}$ and $g_{\theta}$ to HR features $F_{\mathrm{HR}}^{\mathrm{RGB}}$ and $F_{\mathrm{HR}}^{\mathrm{U}}$, respectively
6:     SR RGB $\mu_{\mathrm{HR}}$ and SR uncertainty $\sigma_{\mathrm{HR}}$ from HR features, calculate Eq. (6)
7:     HR depth estimation task:
8:     LR feature $F_{\mathrm{LR}}^{\mathrm{DE}}$ from $I_{\mathrm{LR}}$
9:     LR depth prediction $D_{\mathrm{LR}}$ from $F_{\mathrm{LR}}^{\mathrm{DE}}$, calculate Eq. (9)
10:     Share CIIF parameters $g_{\theta}$
11:     Up-sample $F_{\mathrm{LR}}^{\mathrm{DE}}$ with CIIF of parameters $g_{\theta}$ to HR feature $F_{\mathrm{HR}}^{\mathrm{DE}}$
12:     HR depth prediction $D_{\mathrm{HR}}$ from $F_{\mathrm{HR}}^{\mathrm{DE}}$, calculate Eq. (10)
13:     FD loss:
14:     Transform $F_{\mathrm{HR}}^{\mathrm{RGB}}$ into $F_{\mathrm{HR}}^{\mathrm{T}}$, calculate Eq. (7)
15:     HR Transformed depth $D_{\mathrm{HR}}^{\mathrm{T}}$ from $F_{\mathrm{HR}}^{\mathrm{T}}$, calculate Eq. (8)
16:     Update framework with Eq. (11)
17:  end for
@@
where $N$ is the number of pixels, and $I^{\mathrm{GT}}_{\mathrm{HR}}$ is HR GT image. Except for constraining that the reconstructed high-resolution ODI should be similar to ground truth, $\mathcal{L}_{\mathrm{UISR}}$ additionally emphasizes high-frequency regions that exhibit high uncertainty.